OpenStudy (anonymous):

8. You research the cost of a gallon of gasoline over several years to look for a trend. The table shows your data. What is a line of best fit for the data? Based upon these prices, how much would you expect to pay in 2026?

OpenStudy (anonymous):

Year 1980 1984 1988 1992 1996 2000 2004 2008 2012 Price per Gallon $1.25 $1.27 $1.05 $1.28 $1.43 $1.68 $2.11 $3.72 $3.99

OpenStudy (anonymous):

y = 0.347x + 0.856; $9.88 y = .087x – 171.26; $5.01 y = 0.134x + 1.25; $7.41 y = 0.087x + 0.587; $4.59